6(d+6)=4(d-6)+2d
6d+36=4d-36+2d
6d+4d+2d=12d
12d-36 = -36-36
12d=-72
12/12 and 72/12
d=6 NOT ON MY ANSWER SHEET
a. no solution
b. -60
c.-5
d. all numbers THANK YOU

Answer by ewatrrr(24785)

 
Hi,
6(d+6)=4(d-6)+2d
6d+36=4d-24+2d |Always recommend re-checking work done
6d + 36 = 6d - 24 |subracting 6d from both sides of the equation
36 ≠ -24 There is NO solution